builder bug company allocates overhead at $9 per direct labor hour. Job A45 required 5 boxes of direct materials at a cost of 30 per box and took employees 12 hours to complete. employees earn $15 per hour.
Answer #1
Job Cost has 3 components viz Dir Materail, Dir Labor & OHs

SO Job cost will be :-
1. Dir material = 5Box*$30 per box = $150
2. Dir Lab = 12Hrs*$15 per hr = $180
3. OH = $9 per DL hr*No of DLhrs = $9*12 = $108
------------------------------------------------------------
SO Total Cost of Job A45 is $438
Answer #2
Overheads = 9 * 12 = $108
Direct materials = 5 * 30 = $ 150
Employee costs = 12*15 = $180
So,
Total cost of the job = 108+150+180 = $438
Answer #3
The Total cost of the job is :

(a) Overheads = $9 per labour hour * 12 employee hours = $108
(b) Direct materials = 5 boxes * 30 per box = $ 150
(c) Employee costs = 12*15 = $180

Therefore total cost of the job = a +b + c = 108 + 150 + 180 = $ 438 Answer
Answer #4
total cost =30 * 5 +15 *12

=330
Answer #5
Overheads = 9 * 12 = $108
Direct materials = 5 * 30 = $ 150
Employee costs = 12*15 = $180
So,
Total cost of the job = 108+150+180 = $438
Answer #6
total cost is the summation= a +b + c = 108 + 150 + 180 = 438 dollar